A box has a dimensions of 150cm x 100cm x 120cm and a mass of 260000 cg. Calculate the density of the box in kg/L.
1
Expert's answer
2020-12-10T07:49:40-0500

p=m/V.

m=260000 kg.

V= 150 x 100 x 120 = 1 800 000 cm3 = 1800 L.

p= 260000/1800 = 144,444444 kg/L.